OpenStudy (anonymous):

Two angles are complimentary. Their measures are in a ratio of 8:1.Find their measures using proportions. Show your work.

jimthompson5910 (jim_thompson5910):

"Two angles are complimentary" means that one angle is x and the other is 90-x degrees "Their measures are in a ratio of 8:1" tells us that x/(90-x) = 8/1 Solve for x x/(90-x) = 8/1 x = 8(90-x) x = 720 - 8x x+8x = 720 9x = 720 x = 720/9 x = 80 So x = 80, which means 90-x = 90-80 = 10 Therefore, the two angles are 80 and 10